In a circular piping, water flows with 2 m/s at 298 K and 1 atm. The piping has a diameter of 0.025 m. Determine the mass transfer coefficient and the Schmidt number of water in the air inside with a length of 3 m.
Answer #1

Solution V = 2 m/sec T= 298 K P=1 atm + 9 d = 0.025m 9 L L=3m p = 999 999 yg | mel, M = 1·1*109 PAR = 2.5*105 m² msec sec # R# Sc= 9 M S DAB = 0.044044 Altres # Re = 45409.1 > 4100 so tusbudent flow 1:1¥10 3 % hoc DAB 9994*25*105* derece F Characteri

In case of any doubt or you need any further help regarding this solution please comment below.

Thank you......